An electric current is flowing in a circular wire of radius a. At What distance from the centre on the axis of the circular wire will the magnetic field be 1/8th of it's value at the centre


1
Expert's answer
2021-06-30T10:19:02-0400

"B_1=\frac B8,"

"B_1=\frac{\mu_0}{2}\frac{a^2I}{(a^2+x^2)^{\frac 32}},"

"B=\frac{\mu_0 I}{2a},"

"\frac{a^2}{(a^2+x^2)^{\frac 32}}=\frac1{8a},"

"4a^2=a^2+x^2,"

"3a^2=x^2,"

"x=±a\sqrt 3."